OpenStudy (anonymous):

Solve 4|3w + 1| – 7 = 17

myininaya (myininaya):

You have two equations to solve because |3w+1| can be 3w+1 for when 3w+1>=0 or -(3w+1) for when 3w+1<0 so two equations equation 1 to solve is : 4(3w+1)-7=17 equation 2 to solve is : -4(3w+1)-7=17

OpenStudy (acxbox22):

or...